• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 20:00:41 on June12,2010

    Entry number 326086

    This entry is a followup to: 326083

    keyword=adjust cavity gains

    I found the cavities had a wrong gain.

    The gains were left as per halog 325892 (for 50 nA running)
    So I restored them as per halog 325821 (for 50 uA running)

    I've restored the high current running mode.
    This occurred around event 300K in run 4485.

    I'm not sure what the implications are. At worst we lose some
    redundancy (the cavities are redundant).

    Starting another run (4486) to check cavities.
    But the BCMs still show a sawtooth. Tres bizarre.